Exercise 1.

Starts now and is due in by the end of week seven.

You are asked to read the paper

Pollock and Soffa, April 1992, “Incremental Global Reoptimization of Programs”, ACM Transactions on Programming Languages and Systems Vol 14 No 2

and to provide a summary of its contents in 1000-1500 words, including a discussion of the potential usefulness of this approach compared to modular programming.

This need not be formatted with L^{A}TEX. Clear handwritten or ASCII plain text work is quite acceptable.